# Characterizing Hitchin representations over real closed fields

### v2

@@ 2. Real algebraic geometry
In this section we recall general definitions and results from real algebraic geometry and set up notation. We refer the reader to [BCR98], in particular Chapters 1, 2 and 5, for more details and proofs. Let $\mathbb{F}$ be an ordered field. It is real closed if every positive element is a square and every odd degree polynomial has a root. Examples include the real numbers, the real algebraic numbers and the field of Puiseux series with coefficients in $\mathbb{R}$. Every ordered field has a real closure, i.e. an algebraic field extension that is real closed and that is extending the order. The main objects of study in real algebraic geometry are semi-algebraic sets.
@@
Let $X\subseteq\mathbb{F}^{d}$ and $Y\subseteq\mathbb{F}^{m}$ be two semi-algebraic sets. A map $f\colon X\to Y$ is called semi-algebraic if its graph $\textrm{Graph}(f)\subseteq X\times Y$ is semi-algebraic in $\mathbb{F}^{d+m}$.
@@
Let $X\subseteq\mathbb{F}^{d}$ and $Y\subseteq\mathbb{F}^{m}$ be two semi-algebraic sets, and $f\colon X\to Y$ a semi-algebraic map. Then $(\textrm{Graph}(f))_{\mathbb{K}}$ is the graph of a semi-algebraic map $f_{\mathbb{K}}\colon X_{\mathbb{K}}\to Y_{\mathbb{K}}$, that is called the $\mathbb{K}$-extension of $f$. Furthermore, $f$ is injective (respectively surjective, respectively bijective) if and only if $f_{\mathbb{K}}$ is injective (respectively surjective, respectively bijective), and f is continuous if and only if $f_{\mathbb{K}}$ is continuous.
@@
3. The $\mathbb{F}$-Hitchin component
@@
3.1. The semi-algebraic structure of $\mbox{\rm PSL}(n,\mathbb{F})$
@@
3.2. The semi-algebraic structure of homomorphism spaces
@@
induces a homeomorphism between $\mbox{\rm Hom}(\pi_{1}(S),G)$ and its image $X_{F}(\pi_{1}(S),G)$. Note that $X_{F}(\pi_{1}(S),G)$ is a semi-algebraic subset of some $\mathbb{R}^{d}$. If $F^{\prime}$ is a different choice of a generating set for $\pi_{1}(S)$, then $X_{F}(\pi_{1}(S),G)$ and $X_{F^{\prime}}(\pi_{1}(S),G)$ are semi-algebraically homeomorphic. From now on we drop the choice of a generating set from the notation. Denote by $X^{\textnormal{red}}(\pi_{1}(S),G)$ the image under the map ev of the space of reductive homomorphisms $\mbox{\rm Hom}_{\textnormal{red}}(\pi_{1}(S),G)$. By the theory of Richardson-Slodowy [RS90, Section 4], $X^{\textnormal{red}}(\pi_{1}(S),G)$ is real semi-algebraic as a subset of $X(\pi_{1}(S),G)$. From now on we will identify $\mbox{\rm Hom}(\pi_{1}(S),G)$ and $\mbox{\rm Hom}_{\textnormal{red}}(\pi_{1}(S),G)$ with their respective images under the map ev. Let $\mathbb{F}$ be a real closed field containing $\mathbb{R}$.
@@
3.3. Semi-algebraic models for quotients
Let $G$ be either $\mbox{\rm SL}(n,\mathbb{R})$ or $\mbox{\rm PSL}(n,\mathbb{R})$. The $G$-character variety $\chi(S,G)$ is the topological quotient of the space of reductive homomorphisms $\mbox{\rm Hom}_{\textnormal{red}}(\pi_{1}(S),G)$ modulo $G$-conjugation action. We would like to describe its semi-algebraic structure.
@@ 3.3. Semi-algebraic models for quotients
A semi-algebraic model for the character variety has been given in [RS90, Theorem 7.6].
@@
If $\mathcal{C}\subseteq\mbox{\rm Hom}(\pi_{1}(S),G)$ is a $G$-invariant connected component—hence semi-algebraic by 2.5—consisting of reductive homomorphisms, we can define in an analogous way a semi-algebraic model for the quotient $\mathcal{C}/G$. Indeed, the restriction to $\mathcal{C}$ of the semi-algebraic model for the whole character variety coming from [RS90, Theorem 7.6] provides a semi-algebraic model for $\mathcal{C}/G$. In general we can find different such models, which are related by semi-algebraic homeomorphisms, see 3.6. In fact, for connected components of geometric significance, there are often several semi-algebraic models that exploit their geometric interpretation. For example in Section 7 we introduce a variant of the Bonahon–Dreyer coordinates for the Hitchin component, which generalize the shear coordinates for Teichmüller space (the case of $\mbox{\rm PSL}(2,\mathbb{R})$) developed by Thurston [Thu98, Section 9] and [Bon96, Theorem A].

### v3

@@ 2. Real algebraic geometry
In this section we recall general definitions and results from real algebraic geometry and set up notation. We refer the reader to [BCR98], in particular Chapters 1, 2,5 and 7, for more details and proofs.
2.1. Semi-algebraic sets
Let $\mathbb{F}$ be an ordered field. It is real closed if every positive element is a square and every odd degree polynomial has a root. Examples include the real numbers, the real algebraic numbers and the field of Puiseux series with coefficients in $\mathbb{R}$. Every ordered field $\mathbb{K}$ has a real closure $\overline{\mathbb{K}}^{r}$, i.e. an algebraic field extension that is real closed and that is extending the order. The main objects of study in real algebraic geometry are semi-algebraic sets.
@@ 2.1. Semi-algebraic sets
Let $X\subseteq\mathbb{F}^{d}$ and $Y\subseteq\mathbb{F}^{m}$ be two semi-algebraic sets. A map $f\colon X\to Y$ is called semi-algebraic if its graph $\textnormal{Graph}(f)\subseteq X\times Y$ is semi-algebraic in $\mathbb{F}^{d+m}$.
@@
Let $X\subseteq\mathbb{F}^{d}$ and $Y\subseteq\mathbb{F}^{m}$ be two semi-algebraic sets, and $f\colon X\to Y$ a semi-algebraic map. Then $(\textnormal{Graph}(f))_{\mathbb{K}}$ is the graph of a semi-algebraic map $f_{\mathbb{K}}\colon X_{\mathbb{K}}\to Y_{\mathbb{K}}$, that is called the $\mathbb{K}$-extension of $f$. Furthermore, $f$ is injective (respectively surjective, respectively bijective) if and only if $f_{\mathbb{K}}$ is injective (respectively surjective, respectively bijective), and f is continuous if and only if $f_{\mathbb{K}}$ is continuous.
@@
2.2. Real spectrum of a ring
In this section we introduce the notion of the real spectrum of a ring. We follow [BCR98, Chapters 7.1 and 7.2]. Let $A$ be a commutative ring with $1$. In our examples, $A$ will often be a polynomial ring, compare Section 2.3.
Definition 2.10.
The real spectrum $\textnormal{RSp}(A)$ is the topological space
$\textnormal{RSp}(A)=\{(\mathfrak{p},\leq)\mid\mathfrak{p}\subseteq$ $A\textrm{ prime ideal },$
[…]
together with the following subbasis of open sets: For $a\in A$ let
$\mathcal{U}(a)\coloneqq\{(\mathfrak{p},\leq)\in\textnormal{RSp}(A)\mid\overline{a}^{\mathfrak{p}}>0\},$
where $\overline{a}^{\mathfrak{p}}$ is the image of $a$ in $\textnormal{Frac}(A/\mathfrak{p})$ under the homomorphism
$A\to A/\mathfrak{p}\to\textnormal{Frac}(A/\mathfrak{p}).$
This topology is called the spectral topology on $\textnormal{RSp}(A)$.
Contrary to the spectrum of a ring (where we consider all prime ideals), we restrict our attention here to so called real ideals, which motivates the name of the real spectrum. An ideal is called real if, whenever $a_{1}^{2}+\ldots+a_{k}^{2}\in I$ for some $a_{1},\ldots,a_{k}\in A$, we have $a_{i}\in I$ for all $i=1,\ldots k$, see [BCR98, Definition 4.1.3]. By [BCR98, Lemma 4.1.6] we see that a prime ideal $I\subseteq A$ is real if and only if the fraction field of $A/I$ is orderable.
If $k$ is a field, the real spectrum $\textnormal{RSp}(k)$ of $k$ is homeomorphic to the set of orders on $k$ together with the Harrison topology [BCR98, Example 7.1.4 a)]. It is non-empty if and only if $k$ is orderable. The real spectrum of $\mathbb{Z}$ is one point corresponding to the zero prime ideal $(0)$ and the unique order on the fraction field $\mathbb{Q}$ of $\mathbb{Z}$.
Example 2.11 ([BCR98, Example 7.1.4 (b), 7.5.2]).
Let $\mathbb{R}[X]$ be the polynomial ring in one variable. Since $\mathbb{R}$ is real closed and $\mathbb{R}[X]$ is a principal ideal domain, all non-zero prime ideals are generated by an irreducible polynomial which is either of degree one or two. Irreducible polynomials of degree one correspond to maximal ideals, hence to elements of $\mathbb{R}$, with residue field $\mathbb{R}$. Since $\mathbb{R}$ is real closed it has a unique order. Irreducible polynomials of degree two correspond to algebraic extensions of $\mathbb{R}$ of degree two, which are algebraically closed, and hence not orderable. We also need to describe the orders on the field of rational function $\mathbb{R}(X)$, which is the fraction field of $\mathbb{R}[X]$ corresponding to the prime ideal $(0)$. It suffices to order the variable $X$ with respect to $\mathbb{R}$. More precisely, we define the following orders. For $\lambda\in\mathbb{R}$ we set
  • $\lambda^{+}$: We have $\lambda<_{\lambda^{+}}X$, but $X<_{\lambda^{+}}\mu$ for every $\mu\in\mathbb{R}$ with $\lambda<\mu$.
  • $\lambda^{-}$: We have $X<_{\lambda^{-}}\lambda$, but $\mu<_{\lambda^{-}}X$ for every $\mu\in\mathbb{R}$ with $\mu<\lambda$.
  • +$\infty$+ ∞: We have $\mu<_{+\infty}X$ for all $\mu\in\mathbb{R}$.
  • $\infty$- ∞: We have $X<_{-\infty}\mu$ for all $\mu\in\mathbb{R}$.
It turns out that this list is the complete set of total orders on $\mathbb{R}[X]$. Putting everything together we obtain
$\textnormal{RSp}(\mathbb{R}[X])=\{(\langle X-\lambda\rangle,\leq_{\mathbb{R}})\mid\lambda\in\mathbb{R}\}\cup\{((0),\leq_{o})\mid o\in\{\lambda^{\pm},\pm\infty\}\}.$
There is an equivalent way of viewing points in the real spectrum.
Proposition 2.12 ([BCR98, Proposition 7.1.2]).
Points $(\mathfrak{p},\leq)\in\textnormal{RSp}(A)$ are in bijection with equivalence classes of ring homomorphisms $[\varphi\colon A\to\mathbb{F}]$ for a real closed field $\mathbb{F}$, where we consider the equivalence relation generated by proclaiming two homomorphisms $\varphi\colon A\to\mathbb{F}$ and $\varphi^{\prime}\colon A\to\mathbb{F}^{\prime}$ to be equivalent if there exists an order-preserving field homomorphism $\mathbb{F}\to\mathbb{F}^{\prime}$ such that the diagram commutes:
${A}$${\mathbb{F}}$${\mathbb{F}^{\prime}}$$\scriptstyle{\varphi}$$\scriptstyle{\varphi^{\prime}}$
More precisely, given $(\mathfrak{p},\leq)$ we get a homomorphism from $A$ into the real closure of the residue field $\textnormal{Frac}(A/\mathfrak{p})$ with the order $\leq$ by composing the following maps
$\varphi\colon A\to A/\mathfrak{p}\hookrightarrow\textnormal{Frac}(A/\mathfrak{p})\hookrightarrow\overline{(\textnormal{Frac}(A/\mathfrak{p}),\leq)}^{r}.$
On the other hand, given $[\varphi\colon A\to\mathbb{F}]$ for a real closed field $\mathbb{F}$, take $(\ker(\varphi),\leq)$ with the restriction of the order of $\mathbb{F}$ to $\textnormal{Frac}(A/\ker(\varphi))$. Under this identification we have that for $a\in A$
$\mathcal{U}(a)=\{[\varphi\colon A\to\mathbb{F}]\mid\mathbb{F}\textrm{ real closed},\,\varphi(a)>0\}.$
Theorem 2.13 ([BCR98, Proposition 7.1.12]).
The topological space $\textnormal{RSp}(A)$ with its spectral topology is compact (but not necessarily Hausdorff).
In order to define the real spectrum compactification of semi-algebraic sets, we need the following.
Definition 2.14.
A subset of $\textnormal{RSp}(A)$ is called constructible if it can be obtained as a Boolean combination, i.e. finite unions, finite intersections and complements, from the sets $\mathcal{U}(a)$ defined above.
Proposition 2.15 ([BCR98, Proposition 7.1.25 (ii)]).
Let $C\subseteq\textnormal{RSp}(A)$ be a constructible subset endowed with the subspace topology of the spectral topology. The topological space $C^{\textnormal{cl}}$ of closed points of $C$ is a compact Hausdorff space. In particular $\textnormal{RSp}^{\textnormal{cl}}(A)$, the set of closed points of $\textnormal{RSp}(A)$, is a compact Hausdorff space.
Example 2.16 (continuation of 2.11).
The points $((0),\pm\infty)\in\textnormal{RSp}(\mathbb{R}[X])$ are closed. The points $((0),\lambda^{\pm})\in\textnormal{RSp}(\mathbb{R}[X])$ for $\lambda\in\mathbb{R}$ are not closed. We have $\overline{\{\lambda^{\pm}\}}=\{\lambda^{\pm},\lambda\}$. The space $\textnormal{RSp}^{\textnormal{cl}}(\mathbb{R}[X])$ is homeomorphic to the two point compactification of $\mathbb{R}$, i.e. to the closed interval $[0,1]$.
We have the following characterization of closed points of $\textnormal{RSp}(A)$.
Definition 2.17.
Let $A\subseteq A^{\prime}\subseteq\mathbb{F}$ be two subsets of an ordered field $\mathbb{F}$. We say that $A^{\prime}$ is Archimedean over $A$ if for all $a^{\prime}\in A$ there exists $a\in A$ with $a^{\prime}<a$. An ordered field is Archimedean if it is Archimedean over $\mathbb{N}$.
Proposition 2.18 ([Bru88, Proposition 2.2 (e)]).
A point $(\mathfrak{p},\leq)\in\textnormal{RSp}(A)$ is closed if and only if $\textnormal{Frac}(A/\mathfrak{p})$ is Archimedean over $\varphi(A)$, where
$\varphi\colon A\to A/\mathfrak{p}\hookrightarrow\textnormal{Frac}(A/\mathfrak{p})\hookrightarrow\overline{(\textnormal{Frac}(A/\mathfrak{p}),\leq)}^{r}$
is defined as in 2.12.
2.3. Real spectrum compactification of closed semi-algebraic sets
We now describe the real spectrum compactification of closed semi-algebraic sets. Note that these include the algebraic subsets. Let $A\coloneqq\mathbb{R}[X_{1},\ldots,X_{d}]$ the coordinate ring with coefficients in $\mathbb{R}$. Note that $A$ is naturally an $\mathbb{R}$-algebra, hence contains $\mathbb{R}$.
Lemma 2.19.
Let $\varphi\colon A\to\mathbb{F}$, for $\mathbb{F}$ a real closed field, represent a point in $\textnormal{RSp}(A)$. Then $\mathbb{R}\subseteq\mathbb{F}$ and $\varphi$ is $\mathbb{R}$-linear.
Proof.
The ring homomorphism $\varphi$ restricted to $\mathbb{R}$ is injective (since the only ideals of $\mathbb{R}$ are trivial and $\varphi$ sends one to one), hence $\mathbb{R}\subseteq\mathbb{F}$ via $\varphi$. Furthermore, for $\lambda\in\mathbb{R}$ and $f\in A$, we have $\varphi(\lambda f)=\varphi(\lambda)\varphi(f)$, since $\varphi$ is a ring homomorphism. ∎
Proposition 2.20 ([BCR98, Proposition 7.1.5]).
The map
$\Psi\colon\mathbb{R}^{d}\to\textnormal{RSp}(A),\quad v\mapsto(\langle X_{1}-v_{1},\ldots,X_{d}-v_{d}\rangle,\leq_{\mathbb{R}}),$
where $\leq_{\mathbb{R}}$ is the unique order on $\mathbb{R}$, is injective and induces a homeomorphism from $\mathbb{R}^{d}$ with its Euclidean topology, onto its image in $\textnormal{RSp}(A)$ with its spectral topology.
If $X\subseteq\mathbb{R}^{d}$ is a semi-algebraic set given by a Boolean combination of the basic semi-algebraic sets $\mathcal{B}(f_{i})$ for some $f_{i}\in A$ (2.1), then we define $\widetilde{X}$ to be the constructible subset of $\textnormal{RSp}(A)$ given by the same Boolean combination of the open sets $\mathcal{U}(f_{i})$ (2.10). It turns out that $\widetilde{X}$ is intrinsically defined by the semi-algebraic set $X$ (up to homeomorphism) and does not depend on the ambient space $\mathbb{R}^{d}$ in which $X$ is embedded, see [BCR98, Corollary 7.2.4, Remark 7.2.5].
Theorem 2.21 ([BCR98, Theorem 7.2.3]).
  1. (1)
    There is an isomorphism of Boolean algebras
[…]
    $X$ $\mapsto\widetilde{X},$
    $\Psi^{-1}(\widetilde{X}\cap\Psi(\mathbb{R}^{d}))$ $\mapsfrom\widetilde{X}.$
  2. (2)
    $X$ is closed (resp. open) if and only if $\widetilde{X}$ is closed (resp. open).
Proposition 2.22 ([BCR98, Proposition 7.2.7]).
Let $X$ be a closed (respectively open) semi-algebraic subset of $\mathbb{R}^{d}$. Then $\widetilde{X}$ is the smallest (respectively largest) closed (respectively open) subset of $\textnormal{RSp}(A)$ whose intersection with $\Psi(\mathbb{R}^{d})$ is $\Psi(X)$.
With this at hand we are now ready to define the real spectrum compactification of closed semi-algebraic sets.
Definition 2.23.
Let $X\subseteq\mathbb{R}^{d}$ be a closed semi-algebraic subset. Its real spectrum compactification $\textnormal{RSp}(X)$ is the closure of its image
$X\subseteq\mathbb{R}^{d}\xhookrightarrow{\Psi}\textnormal{RSp}(A).$
The definition of the compactification might à priori depend on the embedding $X\subseteq\mathbb{R}^{d}$. Using 2.22, one easily sees that if $X\subseteq\mathbb{R}^{n}$ be a closed semi-algebraic subset of $\mathbb{R}^{d}$, then $\textnormal{RSp}(X)=\widetilde{X}$. This implies that the real spectrum compactification $\textnormal{RSp}(X)$ is intrinsic to $X$, and does not depend on the embedding $X\subseteq\mathbb{R}^{d}$. Note that in [BIPP23, §2.7], they define the real spectrum compactification of a semi-algebraic set by its associated constructible set, which in the case of closed semi-algebraic sets gives the same definition as here.
From now on let $X\subseteq\mathbb{R}^{d}$ be a closed semi-algebraic set. We saw in 2.16 that in general $X$ is not open in $\textnormal{RSp}(X)$ (not even for algebraic sets). Moreover, $\textnormal{RSp}(X)$ with its spectral topology is in general not Hausdorff (2.13), which is something one often desires when compactifying a Hausdorff topological space. Both of these issues can be resolved when considering the subset of closed points.
Definition 2.24.
Let $\textnormal{RSp}^{\textnormal{cl}}(X)\coloneqq\widetilde{X}^{\textnormal{cl}}$ be the subset of closed points in $\textnormal{RSp}(X)$.
Note that $\textnormal{RSp}^{\textnormal{cl}}(X)=\widetilde{X}\cap\textnormal{RSp}^{\textnormal{cl}}(A)$, since $\widetilde{X}$ is closed by 2.21 (2).
Proposition 2.25 ([BIPP23, Corollary 2.32, Proposition 2.33]).
The closed semi-algebraic set $X$ is dense in $\textnormal{RSp}(X)$, and open and dense in $\textnormal{RSp}^{\textnormal{cl}}(X)$.
${\textnormal{RSp}^{\textnormal{cl}}(X)}$${X}$${\textnormal{RSp}(X)}$open and densedense
Let us collect some important properties of the real spectrum compactification of closed semi-algebraic sets. Recall that $\textnormal{RSp}(X)=\widetilde{X}$.
Proposition 2.26 ([BCR98, Proposition 7.5.1]).
Let $X$ be a semi-algebraic set. Then $X$ is semi-algebraically connected if and only if $\widetilde{X}$ is connected. Furthermore, if $X_{1},\ldots,X_{m}$ are the semi-algebraically connected components of $X$, then $\widetilde{X}_{1},\ldots,\widetilde{X}_{m}$ are the connected components of $\widetilde{X}$.
From this remark we immediately obtain that the real spectrum compactification of a closed connected semi-algebraic set is connected. The following proposition tells us that semi-algebraic maps extend continuously to the compactification.
Proposition 2.27 ([BCR98, Proposition 7.2.8]).
Let $X$ and $Y$ be two semi-algebraic sets and $f\colon X\to Y$ a semi-algebraic map. Then there exists a unique map $\tilde{f}\colon\widetilde{X}\to\widetilde{Y}$ such that for all semi-algebraic subsets $Y^{\prime}\subseteq Y$ we have
$\tilde{f}^{-1}(\widetilde{Y^{\prime}})=\widetilde{f^{-1}(Y^{\prime})}.$
If additionally $f$ is a homeomorphism then so if $\tilde{f}$.
3. $\mathbb{F}$-Hitchin component
@@
3.1. Semi-algebraic structure of $\mbox{\rm PSL}(n,\mathbb{F})$
@@
3.2. Semi-algebraic models for character varieties
@@
induces a homeomorphism between $\mbox{\rm Hom}(\pi_{1}(S),G)$ and its image $X_{F}(\pi_{1}(S),G)$. Note that $X_{F}(\pi_{1}(S),G)$ is a semi-algebraic subset of some $\mathbb{R}^{d}$. If $F^{\prime}$ is a different choice of a generating set for $\pi_{1}(S)$, then $X_{F}(\pi_{1}(S),G)$ and $X_{F^{\prime}}(\pi_{1}(S),G)$ are semi-algebraically homeomorphic. From now on we drop the choice of a generating set from the notation.
A representation $\rho\colon\pi_{1}(S)\to G$ is reductive if, seen as a linear representation on $\mathbb{R}^{m}$, it is completely reducible, i.e. a direct sum of irreducible representations. Similarly, if $\mathbb{F}$ is a real closed field extension and $G_{\mathbb{F}}$ the $\mathbb{F}$-extension of $G$ (2.6), we say that $\rho\colon\Gamma\to G_{\mathbb{F}}$ is reductive if, seen as a linear representation on $\mathbb{F}^{n}$, it is completely reducible. Denote the set of reductive homomorphisms from $\pi_{1}(S)$ to $G$ by $\mbox{\rm Hom}_{\textnormal{red}}(\pi_{1}(S),G)$. Any irreducible representation is reductive. It follows from [Bou12, §20, p. 376 Corollaire a)] that reductive representations are up to conjugation in $\mbox{\rm GL}(n,\mathbb{R})$ determined by their trace function. The same holds true for reductive representations into $\mbox{\rm GL}(n,\mathbb{F})$, where $\mathbb{F}$ is a real closed field.
Denote by $X^{\textnormal{red}}(\pi_{1}(S),G)$ the image under the map ev of the space of reductive homomorphisms $\mbox{\rm Hom}_{\textnormal{red}}(\pi_{1}(S),G)$. By the theory of Richardson-Slodowy [RS90, Section 4], $X^{\textnormal{red}}(\pi_{1}(S),G)$ is real semi-algebraic as a subset of $X(\pi_{1}(S),G)$. From now on we will identify $\mbox{\rm Hom}(\pi_{1}(S),G)$ and $\mbox{\rm Hom}_{\textnormal{red}}(\pi_{1}(S),G)$ with their respective images under the map ev. Let $\mathbb{F}$ be a real closed field containing $\mathbb{R}$.
@@
The group $G$ acts on $\mbox{\rm Hom}(\pi_{1}(S),G)$ by conjugation, i.e. for all $\rho\in\mbox{\rm Hom}(\pi_{1}(S),G)$ and $g\in G$ we have
$(g.\rho)(\gamma)\coloneqq g\rho(\gamma)g^{-1}\,\textrm{ for all }\gamma\in\pi_{1}(S).$
The subset of reductive homomorphisms is invariant under the action of $G$ on $\mbox{\rm Hom}(\pi_{1}(S),G)$ by conjugation, which allows us to define the following.
Definition 3.5.
The character variety is the topological quotient
$\chi(S,G)\coloneqq\mbox{\rm Hom}_{\textnormal{red}}(\pi_{1}(S),G)/G,$
where $G$ acts on $\mbox{\rm Hom}_{\textnormal{red}}(\pi_{1}(S),G)$ by conjugation.
We would like to describe the semi-algebraic structure of $\chi(S,G)$.
@@
A semi-algebraic model for the character variety has been given in [RS90, Theorem 7.6]. For more details on this construction, especially in the context of character varieties, we recommend [BIPP23, §7.2].
@@
If $\mathcal{C}\subseteq\mbox{\rm Hom}(\pi_{1}(S),G)$ is a $G$-invariant connected component—hence semi-algebraic by 2.5—consisting of reductive homomorphisms, we can define in an analogous way a semi-algebraic model for the quotient $\mathcal{C}/G$. Indeed, the restriction to $\mathcal{C}$ of the semi-algebraic model for the whole character variety coming from [RS90, Theorem 7.6] provides a semi-algebraic model for $\mathcal{C}/G$. In general we can find different such models, which are related by semi-algebraic homeomorphisms, see 3.7. In fact, for connected components of geometric significance, there are often several semi-algebraic models that exploit their geometric interpretation. For example in Section 7 we introduce a variant of the Bonahon–Dreyer coordinates for the Hitchin component, which generalize the shear coordinates for Teichmüller space (the case of $\mbox{\rm PSL}(2,\mathbb{R})$) developed by Thurston [Thu22, Section 9] and [Bon96, Theorem A].